A 15.000 g water sample contaminated with mercury (I) is reacted with excess sodium chloride to produce 75.6 mg of precipitate. What was the mass percent of mercury in the sample?
Answer #1

mass of precipitate = 75.6mg = 0.0756

Hg(l) + 2NaCl(aq) ----------> HgCl2(s) + 2Na(s)

1 mole of HgCl2 produced from 1 mole of Hg(l)

271.52g of HgCl2 produced from 200.59g of Hg(l)

0.0756g of HgCl2 produced from = 200.59*0.0756/271.52   = 0.05585g of Hg(l)

mass percent of mercury in the sample   = 0.05585*100/15   = 0.372% >>>>answer